OpenStudy (anonymous):

Solve the system of equations. 3x+5y=-36 y=x-5

OpenStudy (anonymous):

let's replace y=x-5 at the first equation: 3x+5(x-5)=-36 3x+5x-25=-36 8x=-11 x=-11/8 and y=-11/8-5 y=(-11-40)/8 y=-51/8

OpenStudy (anonymous):

@mathsucks_0

OpenStudy (anonymous):

okay so x=-11/8 and y=-51/8... do I need anything else to solve this problem?

OpenStudy (anonymous):

@M_lowreen

OpenStudy (anonymous):

no , so the solution is x=-11/8 and y=-51/8 to make it sure just replace x and y that we found at the equations above @mathsucks_0

OpenStudy (anonymous):

okay thanks!

OpenStudy (anonymous):

you're welcome! pls can you give me a medal? i really need them :) @mathsucks_0

OpenStudy (anonymous):

how do I do that?